### v1

# DeepVoting: Learning Voting Rules with Tailored Embeddings

Abstract
Aggregating the preferences of multiple agents into a collective decision is a common step in many important problems across areas of computer science including information retrieval, reinforcement learning, and recommender systems. As Social Choice Theory has shown, the problem of designing algorithms for aggregation rules with specific properties (axioms) can be difficult, or provably impossible in some cases. Instead of designing algorithms by hand, one can learn aggregation rules, particularly voting rules, from data. However, the prior work in this area has required extremely large models, or been limited by the choice of preference representation, i.e., embedding. We recast the problem of designing a good voting rule into one of learning probabilistic versions of voting rules that output distributions over a set of candidates. Specifically, we use neural networks to learn probabilistic social choice functions from the literature. We show that embeddings of preference profiles derived from the social choice literature allows us to learn existing voting rules more efficiently and scale to larger populations of voters more easily than other work if the embedding is tailored to the learning objective. Moreover, we show that rules learned using embeddings can be tweaked to create novel voting rules with improved axiomatic properties. Namely, we show that existing voting rules require only minor modification to combat a probabilistic version of the No Show Paradox.
1 Introduction
Research in Computational Social Choice (COMSOC) and Algorithmic Game Theory (AGT) focuses heavily on the design and analysis of mechanisms for collective decision making. Canonically, agents arrive with individual preferences over a set of alternatives or outcomes, and a centralized mechanism must aggregate these preferences into a shared choice (voting and selection) or allocation of items (matching and auctions) Shoham and Leyton-Brown (2008). The goal is to design mechanisms with certain desirable properties, characterized by axioms; i.e. optimizing a particular objective or satisfying certain constraints.
One of the central results in social choice is Arrow’s General Impossibility Theorem Arrow et al. (1963), which identifies a set axioms that no collective mechanism can satisfy at once. Following Arrow, decades of research has produced myriad theorems showing which axioms are satisfied by which mechanisms and which sets of axioms lead to an impossibility result Sen (2018). On the algorithmic and computational side this includes properties of optimality and computational complexity Brandt et al. (2016).
Finding rules that satisfy a given set of axioms can be difficult, especially when it is unknown if such a rule exists. Hence, recent work has turned to machine learning techniques to design novel mechanisms. This idea has been applied in areas from auctions to voting rules to matchings Xia (2013); Sandholm (2003); Curry et al. (2022); Ravindranath et al. (2021). Previous work on learning voting rules has been hampered by technical challenges, including extremely large/sophisticated neural nets Anil and Bao (2021), limited data Burka et al. (2022), or failure to account for the full consequences of the design choices Firebanks-Quevedo (2020).
Our approach to improving the learning of existing and novel voting rules is to use hand-crafted embeddings derived from the social choice literature. As we show, these embeddings enable fast learning with fewer parameters and the ability to scale to larger populations of voters. Our embeddings reduce the input size of our neural net, greatly reducing the number of model parameters. Anil and Bao (2021) found that using a multi-layer perceptron (MLP), i.e., a deep neural network, to learn voting rules was limited by the network’s fixed input size, and so they introduced more sophisticated architectures to permit scaling and pad smaller profiles. In their tests to see which architectures scaled best with the number of voters, their MLP architecture was necessarily excluded. By contrast, our embeddings enable the size of the input layer of the MLP to be independent of the number of voters, thereby enabling greater generalization with scale. Similarly, when we train or test our model with preference profiles containing fewer voters, our input does not require any form of padding. As we show, when using embeddings to learn particular rules or axioms, the choice of embedding must correspond to the learning objective.
Sen (2018) observed that mechanisms can be examined in terms of what information they use and ignore from preference profiles. A key contribution of our work is a more complete understanding of the relationship between the choice of embedding and the resulting learnability and efficiency of the mechanisms. We observe that, like any compression algorithm, embeddings can be lossy and impose a bound on the learnability of rules and axioms. Indeed, we show that certain embeddings lose the required information needed to learn particular rules. Our experiments have also lead us to pose new theoretical questions about the information preservation of embeddings.
When most people think of voting they think of classical deterministic rules that take in preferences over a small set of candidates and return a single winner Zwicker (2016); Taylor (2005). However, the full space of social choice mechanisms is much richer with mechanisms varying by the data types of their inputs and outputs. For example, voters may give approval ballots, rankings, or weightings to different candidates; the outcome of the mechanism may be a single winning candidate, collection of winners, or ordering of the candidates, among other options.
We study probabilistic social choice functions (PSCFs) which take a profile over candidates and return a lottery (probability distribution) over the candidates Brandt (2017). PSCFs offer several advantages when learning voting rules with neural networks as compared to single-winner voting rules. First, since we are outputting a lottery over the alternatives, we are not as concerned with tie-breaking. Tie-breaking introduces complications into the design and analysis of voting rules Aziz et al. (2013b), and random tie-breaking is not learnable. Second, PSCFs provide a natural connection between the discrete formulations of rules and axioms and the construction of continuous loss functions for training based on divergences between distributions, e.g. L1.
After first showing that existing rules can be efficiently learned using tailored embeddings, we address the challenge of taking existing rules and tweaking them to improve their axiomatic properties. In particular, we focus on the No Show Paradox in which a voter can induce an outcome they prefer by abstaining instead of voting Moulin (1988). Certain single-winner rules, like Plurality, Borda, and Simpson-Kramer are known to satisfy the Participation axiom, so no voter can be made better off by abstaining and the paradox does not arise. However, most other common voting rules exhibit this paradox. In our experiments, we take models trained to learn existing PSCFs and retrain them using a loss function that adds in a continuous relaxation of the Participation axiom and show which rules can be adjusted to be more resistant to the paradox without sacrificing accuracy.
We choose the Participation axiom in particular because it is an inter-profile axiom, which requires reasoning about counterfactuals on what the preference profile could have otherwise been had the voters behaved differently. Inter-profile axioms are particularly challenging for learning from data because they increase the computational complexity of computing loss functions. It also requires that the model be able to take profiles of different sizes (differing by one voter) as input, which our embeddings enable us to do. Since abstention can be a strategic behavior by voters, learning the Participation axiom is closely related to Automated Mechanism Design, which focuses on creating desirable economic mechanisms for strategic agents and “shifts the burden of design from man to machine” Sandholm (2003).
2 Related Work
Xia (2013) and Procaccia et al. (2009) proposed incorporating voting axioms into a machine-learning framework as a means of evaluating learned social choice mechanisms. In the space of auction design and matching there has been work on using neural nets for better mechanisms Dütting et al. (2019); Pavlov (2011); Malakhov and Vohra (2008) including learning new types of auction mechanisms Curry et al. (2022) as well as complex preference structures Peri et al. (2021). More recently, the work of Ravindranath et al. (2021) has looked at how to learn new allocation mechanisms that bridge the gap between stability (as compared to the deferred acceptance algorithm Gale and Shapley (1962)) and strategyproofness (as compared to random serial dictatorship (RSD) Aziz et al. (2013a)). While the work of Ravindranath et al. (2021), Firebanks-Quevedo (2020), and most recently Anil and Bao (2021) has shown promise for learning mechanisms, these efforts do not closely consider the role of embeddings.
While formal proposals to learn voting rules date back over a decade Xia (2013), considerable attention to learning voting rules has increased in recent years. Kujawska et al. (2020) and Burka et al. (2022) used several common machine learning methods to mimic existing voting rules. However, both of these works overlooked the importance of the choice of embedding in the role of learning, finding that certain rules were “easier” to learn. Subsequently, Anil and Bao (2021) showed that PIN architectures have the advantage of better generalization to larger numbers of voters. We build on this work by showing that we can achieve high accuracy efficiently with small MLPs by using hand-crafted embeddings.
Procaccia et al. (2009) showed that positional scoring rules are efficiently PAC learnable, but learning pairwise comparison-based voting rules requires an exponential number of samples. While we do not escape the asymptotic limits, we examine two embeddings based on tournament graphs that are design to facilitate more efficient learning of pairwise-comparison based rules. Firebanks-Quevedo (2020) use one measure of optimality (Condorcet consistency) and strategyproofness for learning. However, the manipulations of the chosen embedding cannot learn strategyproofness, leading to poor results. Finally, Wilson (2019) focus on the problem of learning a voting rule given a set of pair-wise relations and properties that must hold for the optimization criteria. However, this work is focused on the possibility of learning these functions and does not employ any machine learning techniques.
The loss function chosen by Armstrong and Larson (2019) was a function of the profile and outcome, and thus could learn a rule but not inter-profile axioms like Participation. Recently, Mohsin et al. (2022) focused on the problem of designing and/or learning fair and private rules, proving that under measures of differential privacy there is an upper bound on the trade-off between group fairness and efficiency.
Learning voting rules bears some similarity to the well studied area of learning to rank (L2R) from the machine learning literature Cao et al. (2007). In L2R one typically is concerned only with accurate recovery of the population preference and not the axioms or properties of the aggregation method itself (e.g., fairness). Indeed, one can think of our work enforcing inter-profile axioms on the learned aggregation procedures as an important step.
3 Preliminaries
Agents and Preference Profiles
Let $V$ be a set of $n$ voters and $C$ a set of $m$ candidates. Each voter $i\in V$ reports a strict order $x_{i}$ over all the candidates in $C$ as their ballot. We will denote that $i$ strictly prefers $a$ over $b$ by $a\succ_{i}b$ for $a,b\in C$. There are $m!$ possible ballots, or ways to strictly order (permute) the candidates in $C$. A list of $n$ ballots, one for each voter, constitutes a profile $X=(x_{i})_{i\in V}$. Voter $i$ ranks candidate $a$ at position $x_{i}(a)\in[m]$, using $[k]=\{1,\ldots,k\}$. Let $\mathcal{X}$ be the set of all possible profiles.
Probabilistic Social Choice Functions
A probabilistic social choice function (PSCF) is a function $f:\mathcal{X}\rightarrow\Delta(C)$ that takes a profile $X\in\mathcal{X}$ as input and returns a lottery, or probability distribution $f(X)\in\Delta(C)$ over the set of candidates in the profile, where $\Delta(C)$ is the set of all lotteries over $C$. Let $\mathcal{F}$ be the set of all such PSCFs.
Any PSCF can be used to construct a non-deterministic voting rule by sampling a winner from the lottery. A PSCF that places all probability mass on a single candidate for all profiles is deterministic. If the candidate that receives all of the probability mass is the same candidate for all profiles, then it is a dictatorial voting rule. Many PSCFs we consider return for all profiles a lottery that is a uniform distribution over a non-empty subset of the candidates, i.e., there are multiple potential winners that we would have to choose from for a single-winner voting rule. Therefore, let $U(Y)$ denote the uniform distribution over any finite set $Y$. When referring to lotteries over candidates, we let $U(Y)$ denote the distribution that is uniform over $Y\subseteq C$ and zero on $C\backslash Y$.
Embedding
Simple feed-forward neural networks require a fixed-size input for learning and inference, corresponding to the size of their input layer. If we were to learn voting rules using neural networks that take the entire profile as input, then not only does the input layer need to be large $(m\times n)$, but it also prevents scaling up as the number of voters grows. Similarly, if the number of voters shrinks, then the profile would have to be padded carefully in a way that doesn’t negatively impact the model. If we want to learn rules that are agnostic to the number of voters, we need to construct embeddings of fixed-size that retain the relevant information for profiles with any number of voters. Naturally, different embeddings preserve different information from the original profile, leading them to different efficacy when learning different rules and axioms. Note that most rules and axioms in the literature are defined for any positive number of voters, so we would like our learned mechanisms to be similarly agnostic.
Figure 1: Each of our three embeddings derived from a ballot profile. Note that the size of the profile (a) grows in $O(mn)$, while each of our embeddings grows with $O(m^{2})$, which is far smaller when $m<<n$. However, our embeddings do not always preserve all of the information in the original profile.
An embedding $T$ is a function $T:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ mapping profiles to some codomain $\mathcal{X}^{\prime}$. The embeddings we are concerned with are many-to-one mappings. This means multiple different profiles may have the same embedding, i.e. $T(X)=T(\tilde{X})$ for some $X,\tilde{X}\in\mathcal{X}$ where $X\neq\tilde{X}$. In other words, $T$ will not be reversible, and $T(X)$ will not preserve all information about $X$. We denote by $\mathcal{F}^{\prime}$ the set of all probabilistic functions of the form $f^{\prime}:\mathcal{X}^{\prime}\rightarrow\Delta(C)$. Note that while we designate $\mathcal{X}$ to always contain strict orders over candidates, the structure of $\mathcal{X}^{\prime}$ will be different for different embeddings. Our three embeddings are drawn from the voting literature, but are not commonly recognized as embeddings in the machine learning literature.
Definition 1 (Tournament Embedding).
Given a profile, the tournament embedding $T_{T}$ produces a $m\times m$ matrix $M$ where $M[j,k]=1$ if a majority of voters prefer $j\succ_{i}k$, $M[j,k]=0$ if a majority prefer $k\succ_{i}j$, and $M[j,k]=\frac{1}{2}$ if an equal number of voters prefer each candidate (when $n$ is even), for candidate pairs $j,k\in C$.
Definition 2 (Weighted Tournament Embedding).
Given a profile, the weighted tournament embedding $T_{WT}$ produces a $m\times m$ matrix $M$ where $M[j,k]=|\{i\in V:j\succ_{i}k\}|$ for $j,k\in C$.
Observe that $T_{WT}$ contains strictly more information about the original profile than $T_{T}$ as the tournament can be computed from the weighted tournament.
Definition 3 (Rank Frequency Embedding).
Given a profile, the rank frequency embedding $T_{RF}$ produces a $m\times m$ matrix $M$ showing how many voters rank each candidate $c\in C$ in each position $k\in[m]$ where $M[c,k]=|\{i\in V\text{ s.t. }\succ_{i}^{c}=k\}|$.
There is a tension in the literature between rules that use positional information, like scoring rules, and those that rely on majoritarian or pairwise comparison information, like tournament rules Brandt et al. (2014). Note $T_{RF}$ maintains positional information while $T_{WT}$ and $T_{T}$ maintain majoritarian.
Probabilistic Social Choice Functions
We now define our PSCFs. Where necessary, we always break ties lexicographically. Definitions for IRV and Black’s rule are in Appendix A. We include two rules that are typically classified as scoring rules in the voting literature, Borda and Plurality. The outcome of any scoring rule can be exactly computed from $T_{RF}$.
Definition 4 (Borda).
The Borda score of candidate $c\in C$ from profile $X$ is $B(c)=\sum\limits_{i\in V}(m-x_{i}(c))$. Let $W(X)=\operatorname*{arg\,max}\limits_{c\in C}B(c)$ be the subset of candidates with maximum Borda score. The probabilistic Borda rule returns the lottery $U(W(X))$. [^1] Referred to as Borda Max by Endriss (2017).
Definition 5 (Plurality).
The Plurality score of candidate $c\in C$ from profile $X$ is $L(c)=|\{i\in V:x_{i}(c)=1\}|$. Let $W(X)=\operatorname*{arg\,max}\limits_{c\in C}L(c)$ be the subset of candidates with maximum Plurality score. The probabilistic Plurality rule returns the lottery $U(W(X))$.
The rest of our rules are not scoring rules. Copeland is typically classified as tournament rule since it’s outcome can be computed directly from $T_{T}$ Brandt et al. (2014).
Definition 6 (Copeland).
The Copeland score of candidate $c\in C$ from profile $X$ is the number of other candidates it beats in pairwise competition plus $\frac{1}{2}$ times the number of other candidates it ties with in direct competition (if $n$ is even). Let $W(X)$ be the subset of candidates with maximum Copeland score on profile $X$. The probabilistic Copeland rule returns the lottery $U(W(X))$.
The Simpson-Kramer (Maximin) and Schulze rules are each computed from $T_{WT}$. We will call these weighted-tournament rules. Let $G_{X}(C,E)$ be the directed tournament graph with edges corresponding to all positive values of the tournament matrix induced by $X$. Let each directed edge $(a,b)\in E$ have weight $d(a,b)=|i\in V:a\succ_{i}b|$.
Definition 7 (Simpson-Kramer).
Let $W$ be the subset of candidates whose maximum weight incoming edge is minimal in $G_{X}$. The probabilistic Simpson-Kramer work returns the lottery $U(W(X))$.
Definition 8 (Schulze).
For each path from $a$ to $b$ in $G_{X}$, we let the strength of the path be the minimum weight edge in that path. For each pair of candidates $a,b\in C$ with a path from $a$ to $b$, we let $p(a,b)$ be the maximum strength of any path from $a$ to $b$, and let $p(a,b)=0$ otherwise. Finally, let $W(X)=\{a\in C:p(a,b)\geq p(b,a)\text{ for all }b\in C\}$. The probabilistic Schulze rule returns the lottery $U(W(X))$.
Some, but not all, of the rules listed above are Condorcet-consistent, meaning that they place all probability mass on the Condorcet winner whenever one exists. A Condorcet winner is a candidate who beats all other candidates in pairwise competition, which can be inferred from $T_{T}$ or $T_{WT}$.
$T_{RF}$ $T_{WT}$ $T_{T}$
Plurality $\times$ $\times$
Borda ? $\times$
Copeland $\times$
Schulze $\times$ $\times$
Ranked Pairs $\times$ $\times$
Simpson-Kramer $\times$ $\times$
IRV $\times$ ? $\times$
Black’s Rule $\times$ ? $\times$
Table 1: PSFC preservation under embedding
4 PSCF Preservation Under Embedding
We are concerned with what information is preserved by embeddings, and whether this information is sufficient to implement PSCFs, i.e. to learn them perfectly.
Definition 9 (PSCF Preservation).
A PSCF $f:\mathcal{X}\rightarrow\Delta(C)$ is preserved by embedding $T:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ if $\exists f^{\prime}:\mathcal{X}^{\prime}\rightarrow\Delta(C)$ such that $f^{\prime}(T(X))=f(X)$ for all profiles $X\in\mathcal{X}$.
Proposition 1 says that for an embedding $T$ to preserve a PSCF, there cannot be two profiles with the same embedding under $T$ for which the PSCF returns different lotteries.
Proposition 1.
Embedding $T$ preserves PSCF $f$ if and only if for all pairs of profiles $X,\hat{X}\in\mathcal{X}$ we have $T(X)=T(\hat{X})\Rightarrow f(X)=f(\hat{X})$.
Some embeddings preserve strictly greater information than others. For example, lexicographically sorting the preference orders in a profile preserves all information necessary to compute $T_{WT}$, and $T_{WT}$ preserves all information necessary to compute $T_{T}$ from a profile. This implies that if $T_{T}$ preserves a function $f$, then $T_{WT}$ must preserve $f$ as well.
Proposition 2.
Suppose that for $T:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ there exist $T_{1}:\mathcal{X}\rightarrow\hat{\mathcal{X}}$ and $T_{2}:\hat{\mathcal{X}}\rightarrow\mathcal{X}^{\prime}$ such that $T(X)=T_{2}(T_{1}(X))$ for all $X\in\mathcal{X}$. Then for all $f\in\mathcal{F}$, $T$ preserves $f$ only if $T_{1}$ preserves $f$.
Since we can compute $T_{T}$ from $T_{WT}$, $T_{T}$ can only preserve a PSCF if $T_{WT}$ does as well. However, the reverse does not hold. $T_{WT}$ may preserve PSCFs that are not preserved by $T_{T}$. If an embedding preserves a PSCF, then the PSCF is perfectly learnable from the embedding.
Table 1 shows which PSCFs are preserved by our embeddings. See Appendix B for proofs of the negative results in Table 1 where PSCFs are not preserved. The question marks in the table represent rule-embedding pairs for which we could not find such a counterexample from brute force search, and yet do not have a proof that preservation holds.
Figure 2: L1 rule loss on validation set of random profiles for Plurality, Borda, Copeland, and Schulze per epoch.
5 Learning Lotteries from PSCFs
In our first set of experiments, we show that with proper embeddings we can learn PSCFs that generalize common voting rules using network architectures with few parameters. We train each of our 21 rule-embedding pairs separately to compare their performance, and show how the choice of embedding must correspond to the choice of rule.
Experimental Setup
We train our PSCFs on profiles with $n=29$ voters and $m=7$ candidates. For all experiments, profiles are generated uniformly at random, reflecting the impartial culture assumption Black et al. (1958). Like Firebanks-Quevedo (2020), we use the Whalrus package to implement our voting rules. [^2] https://pypi.org/project/whalrus/
Embeddings afford three key advantages: (1) They reduce the size of the input layer of our neural net, which is a fully connected layer, and therefore greatly reduce the number of model weights. All three of our embeddings compress the $n\times m$ profile to an $m\times m$ matrix representation, so the same MLP architecture can be used for all training runs. (2) When a rule is paired with an appropriate embedding, the embedding preserves all the information necessary to learn the rule and removes unnecessary information. (1) and (2) mean that we learn PSCFs faster and more accurately than previous work. (3) Input size no longer depends on the number of voters, which lends itself to better scaling. For $T_{RF}$ and $T_{WT}$ we normalize the input by dividing all elements in the embedding by $n$, e.g. the elements of the $T_{WT}$ now represent the fraction of voters who prefer one candidate to another $\frac{d(a,b)}{n}$ for $a,b\in C$.
We emulate the MLP architecture of Anil and Bao (2021), with 5 fully-connected layers of 120 nodes, ReLU activation functions, and a Softmax layer for the output. The key difference is that our network takes in embedded profiles so the size of our input layer is $m^{2}$ compared to their $nm^{2}$. This brings our total number of model parameters down to just under 50K vs. 200K. We train our models on a set of 48,000 randomly sampled profiles in batches of size 32 for 1000 epochs, for a total of 1.5M gradient steps. The use of embeddings also allows us to test our MLP model on larger profiles without increasing the size of the network, which Anil and Bao (2021) were unable to do for their MLP model, and only provided for their more sophisticated architectures. We trained each model on NVIDIA Volta V100 GPUs using PyTorch, with each run taking $\approx$1 hour. We used the Adam optimizer for each run with an initial learning rate of 0.001, tuning on plateau (patience = 10, factor = 0.5, min_lr = 0.01). We refer to the L1 distance between our model output and the PSCF lottery on a profile as the rule loss. Rule losses presented in Table 2 and Table 3 are from a test set of 10,000 random profiles sampled independently of the training data. All of these models were trained to minimize rule loss for their respective PSCFs.
Learned PSCFs
$T_{RF}$ $T_{WT}$ $T_{T}$
Plurality 0.000706 0.154674 0.150783
Borda 0.002114 0.001468 0.065766
Copeland 0.070568 0.058405 0.000737
Schulze 0.071135 0.060596 0.044539
Simpson-Kramer 0.072062 0.056812 0.045097
IRV 0.092131 0.109611 0.079310
Black’s Rule 0.029726 0.029520 0.037891
Table 2: Average rule loss testing models trained to learn PSCFs on uniformly random profiles $(m=7,n=29)$.
Figure 2 shows plots of our validation losses during training, using data not used during training, to illustrate our observations. The plots for the remaining rules can be found in Appendix C.
The scoring rules, Plurality and Borda, learn very rapidly using $T_{RF}$ with rule losses converging to zero in a small number of epochs, as expected because $T_{RF}$ preserves all scoring rules. For plurality, we see a non-zero plateau for $T_{WT}$ and $T_{T}$, as they do not preserve all information needed to learn the rule, and so there is a non-zero lower bound to the error rate.
One of our most surprising results is that Borda appears to learn just as well from $T_{WT}$ as from $T_{RF}$. This raises the question of whether $T_{WT}$ preserves Borda. To our knowledge, there is no known explicit algorithm for computing the Borda outcome from a weighted tournament. We pose the same question for IRV and Black’s Rule with $T_{WT}$. In Social Choice there is thought to be a tension between positional information, such as that contained in $T_{RF}$, and pairwise comparisons. For example, no scoring rule can be Condorcet-consistent. So if Borda can be computed from weighted tournament graphs, as suggested by our learning experiments, this would be a surprising result. Our search for counterexamples with small numbers of voters and candidates did not yield any results, but this search was limited by computational intractability.
Challenge.
Does the weighted tournament embedding preserves Borda, IRV, or Black’s rule?
As expected, the Copeland rule learns rapidly with the $T_{T}$, converging to near zero loss quickly as Copeland can be computed from $T_{T}$. While any rule that can be computed from $T_{T}$ can also be computed from $T_{WT}$, what we see is that the Copeland rule loss converges to zero much more slowly for the $T_{WT}$, failing to reach the same loss as $T_{T}$ in our experiments after 1000 epochs. We make a similar observation for the Schulze rule. However, unlike Copeland, Schulze can be computed exactly from $T_{WT}$ but not $T_{T}$. This is why we see the loss with $T_{T}$ plateau at a nonzero value for Schulze in Figure 2. And yet, the rule loss for Copeland with $T_{WT}$ fails to surpass $T_{T}$ after 1000 epochs. This highlights both the benefits of choosing the right embedding for the rule, and that embeddings containing more information may lead to slower learning. Similar observation can be made for the Simpson-Kramer rule. For several of our rules, when using $T_{WT}$, the rule does not appear to finish learning after 1000 epochs. Ultimately, no single embedding performs best across all rules.
$T_{RF}$ $T_{WT}$ $T_{T}$
Plurality 0.033898 0.173783 0.188456
Borda 0.016891 0.014817 0.081727
Copeland 0.085109 0.085285 0.000763
Schulze 0.089108 0.100582 0.149869
Simpson-Kramer 0.089382 0.100518 0.054845
IRV 0.114596 0.121245 0.094002
Black’s Rule 0.042577 0.040671 0.206753
Table 3: Rule loss on models trained on random profiles $(m=7,n=29)$ tested on profiles with $n=199$ voters.
Scaling With Number of Voters
Our embeddings give us the ability to work with profiles with different numbers of voters. Although we trained our models only with profiles with 29 voters, we can test with larger and smaller numbers of voters to check generalizability. We test first on profiles with 199 votes (Table 3), and then again with only 9 voters (Table 6 in Appendix D). We see that there is an increase in the loss in relative terms for all rules, but most losses remain quite low compared to their initial losses during training. We note that the Copeland rule with $T_{T}$ scaled the best, with low error in absolute terms across all treatments. All of the rules besides Plurality and Borda generalized better in relative terms with $T_{RF}$ than did Plurality and Borda, however Plurality and Borda maintained the lowest errors with that embedding. The rule-embedding pairs with the highest errors initially, also tended to scale quite well in relative terms. And all rules performed better with 9 voters than with 199 for all embeddings. When scaling to 199 voters, Schulze and Black’s rule did exceptionally poorly with $T_{T}$. Once again, this highlights the importance of choosing an embedding that fits the rule, corresponding to the learning objective. No single embedding does better than the others for scaling across the board.
Figure 3: Validation losses for models trained to learn PSCFs with $m=7,n=29$ and retrained to learn Participation with combined loss on profiles with $m=7,n=9$.
6 Resisting the No Show Paradox
PSCFs based on voting rules can be vulnerable to the No Show Paradox, where a voter prefers the outcome yielded by a rule when they do not vote. The voter therefore has an incentive to abstain rather than report their true preference. A rule for which this cannot occur is said to satisfy the Participation axiom. We now employ transfer learning, taking our already trained models from Section 5 and retraining them with a loss function that adds a term for Participation loss, which is our relaxation of the binary participation axiom to a continuous loss function.
For all definitions below, let $P_{X}$ be a probability distribution derived from profile $X$ by some PSCF $f$ (implicit), and let $P_{X}(c)$ be the probability assigned to candidate $c$. Where the specific profile is not relevant, we will denote simply by $P(c)$ the probability assigned to candidate $c$ by a lottery $P$. We now provide a loss function that measures how well a PSCF satisfies participation or how close a PSCF comes to satisfying participation. We use stochastic dominance to model a voter’s preference between two lotteries based on their preference order over candidates.
Definition 10 (Stochastic Dominance).
Let $\sigma$ be an ordering (or permutation) over the set of candidates $C$, and let $\sigma[k]$ be the $k^{th}$ element of $\sigma$ for $k\in[m]$. Given two lotteries $P$ and $Q$ over $C$, $P$ stochastically dominates $Q$ with respect to $\sigma$ if for all $k\in[m]$, $\sum\limits_{l\leq k}P(\sigma[l])\geq\sum\limits_{l\leq k}Q(\sigma[l])$.
We say that a voter’s abstention leads to an outcome ($P$) they prefer if the new outcome stochastically dominates the outcome ($Q$) that would derive from the true profile, with respect to the voter’s ordering of the candidates $\sigma=x_{i}$. We want our PSCF to be strategyproof with respect to the limited class of strategic abstentions.
Definition 11 (Participation).
A PSCF $f$ obeys participation if, for all profiles, every voter prefers the outcome under $f$ when they vote their true preference to the outcome under $f$ when they abstain (i.e. removed). We say that a voter prefers the outcome $Q$ from voting truthfully over the lottery $P$ from abstaining if $Q$ stochastically dominates $P$.
Since Participation is a binary condition for a PSCF, to learn PSCFs that resist the No Show Paradox, we define a non-binary loss function for Participation based on stochastic dominance.
Definition 12 (Stochastic Dominance Loss).
Given ordering $\sigma$ over $C$, a lottery $P$, and a reference lottery $Q$, we say that the stochastic dominance loss is zero if $P$ stochastically dominates $Q$. If $P$ does not stochastically dominate the reference lottery $Q$, then the loss is equal to $L(P|\sigma,Q)=\max\limits_{k\in[m]}(\sum\limits_{l\leq k}Q(\sigma[l])-\sum\limits_{l\leq k}P(\sigma[l]))$, i.e. the largest difference between the sums of prefixes of the lotteries over all prefixes when the distributions’ supports are ordered by $\sigma$.
Definition 13 (Participation Loss).
Given a profile $X$, Let $P^{i}_{X}$ be the lottery under $f$ when voter $i$ abstains and all others vote truthfully, and let $Q_{X}$ be the lottery under $f$ when voting truthfully. $L(f,X)=\max\limits_{i\in V}L(P^{i}_{X}|\sigma,Q_{X})$
$T_{RF}$ $T_{WT}$ $T_{T}$
Plurality 0.282162 0.412889 0.411734
Borda 0.409378 0.408852 0.423503
Copeland 0.399701 0.380584 0.389514
Schulze 0.413933 0.388911 0.395796
Simpson-Kramer 0.403704 0.395422 0.393872
IRV 0.409915 0.413539 0.426163
Black’s Rule 0.416672 0.415060 0.409416
Table 4: Participation Loss of models trained to learn PSCFs on profiles of size ($m=7$, $n=29$), tested on 1000 random profiles of size ($m=7$, $n=9$).
The results in Table 4 are surprising. The single-winner versions of Borda, Plurality, and Simpson-Kramer do not suffer from the No Show Paradox, and yet our learned models for the PSCFs based on these rules show similar Participation losses to our other rules. Plurality showed a lower initial Participation loss when trained with $T_{RF}$, but not drastically lower. This may be due to our choice of the maximum stochastic dominance loss over all voters for a profile rather than, say, the sum of such losses across all voters. The rest of our rules are known to suffer from the paradox Pérez (2001), although the paradox does not arise frequently Brandt et al. (2019). And it is known that Condorcet-consistency is incompatible with Participation when there are at least 4 candidates and at least 12 voters Brandt et al. (2017). By our measure, almost all learned rules have similar initial Participation loss in Table 4.
Experimental Setup
Retraining uses the same architectures and setup as the initial training. However, we retrain on 900 random profiles with 7 candidates and 9 votes as the No Show Paradox is more likely to occur with fewer voters. Our ability to change the numbers of voters, without padding the profile, is a benefit of our embeddings. To compute our loss, we add Participation Loss and the original L1 rule loss for each profile. We retrain for 200 total epochs per model.
Using fewer voters for training is also more computationally efficient, which is important because computing losses based on $n$ alternative profiles for each profile increases the runtime by $O(n)$. This is a major challenge for all inter-profile axioms that involve counterfactual comparisons. The added complexity hinges on how many different profiles must be considered to evaluate whether an axiom is satisfied Schmidtlein (2022); Schmidtlein and Endriss (2023)
We did not retrain all of our models for Table 5. Instead, for each rule, we retrained models using (1) the embedding(s) known to preserve the rule, if any, and (2) whichever embedding(s) performed best in our initial experiment. In Figure 3 we plot the epoch losses from an independent validation set with 100 disjoint random profiles. Each plot shows the Participation loss, rule loss (L1), and the sum of the two used as the combined loss function for training. Additional results are in Appendix E.
Participation-Adjusted PSCFs
The three rules we trained and retrained using $T_{RF}$ – Plurality, Borda, and Black’s Rule – all learned to minimize their Participation loss efficiently, with little or no increase in the rule loss, as shown in Table 5. Plurality learned most efficiently, being the only experiment in which the axiom loss dropped below the rule loss within 200 epochs. Despite Borda having higher Participation loss initially than we might have expected due to single-winner Borda satisfying Participation, it learned to evade the No Show paradox quite easily. Among the rules trained using $T_{T}$ and $T_{WT}$, Simpson-Kramer minimized the Participation loss most efficiently with $T_{T}$, which is not too surprising as single-winner Simpson-Kramer satisfies participation. However, the rule loss does not converge to zero at the same rate, even with $T_{WT}$. It is surprising that learning is so much poorer for Simpson-Kramer with $T_{WT}$ than $T_{T}$. The other rules retrained with $T_{T}$ and $T_{WT}$ struggled to learn to satisfy Participation without increasing rule loss within 200 epochs. This may be because appearances of the No Show Paradox are quite rare, occurring in about 4% of profiles Brandt et al. (2019).
Rule Embedding Participation Loss Rule Loss
Plurality $T_{RF}$ 0.078363 0.054809
Borda $T_{RF}$ 0.254161 0.029481
Borda $T_{WT}$ 0.239112 0.035640
Borda $T_{T}$ 0.432263 0.061900
Copeland $T_{T}$ 0.127699 0.039600
Schulze $T_{WT}$ 0.175672 0.089309
Schulze $T_{T}$ 0.065567 0.058519
Simpson-Kramer $T_{WT}$ 0.136363 0.105944
Simpson-Kramer $T_{T}$ 0.019293 0.062169
IRV $T_{WT}$ 0.165605 0.113929
IRV $T_{T}$ 0.229106 0.093015
Black’s Rule $T_{RF}$ 0.193693 0.039200
Black’s Rule $T_{WT}$ 0.215214 0.066647
Table 5: Participation Loss and Rule Loss (L1) of selected models retrained on profiles of size ($m=7$, $n=9$), averaged over 160 test profiles.
7 Conclusions and Future Work
In this paper we have shown that not only can we efficiently learn known PSCFs from preference data, but also that we can modify these rules in order to improve them in ways that, to date, have not be possible through traditional algorithmic design methods. We have highlighted the importance of the choice of embedding on the efficiency and quality of the learned rules, finding some surprising results including that rules like IRV, Borda, and Black’s Rule, which are not majoritarian rules, can be learned well from $T_{WT}$. It remains to be seen whether other embeddings can be designed, of size $m\times m$ or smaller, that outperform the embeddings we took from the social choice literature. Different embeddings may be beneficial in particular for rules whose outcomes are NP-Hard to compute or other common axioms.
Acknowledgments
This work was funded in part by GoodLynx and IBM. This material is based upon work supported by the National Science Foundation under Grant #2030859 to the Computing Research Association for the CIFellows Project. Ben Abramowitz was supported by the CIFellows Project and GoodLynx Inc. Nicholas Mattei was supported in part by NSF Awards IIS-RI-2339880, IIS-RI-2007955, IIS-III-2107505, and IIS-RI-2134857. The authors would like to thank Liam Healy for early tinkering and productive discussions and explorations around this topic.
Appendix A Additional Voting Rules
In this section we give full definitions of other voting rules we study.
Instant Runoff Voting is not a scoring rule, but is defined by iteratively using plurality scores.
Definition 14 (Instant Runoff Voting (IRV)).
IRV is a deterministic, iterative voting rule that, in each of $m-1$ rounds, eliminates the candidate with the lowest plurality score and removes them from the preference orders of all voters before the next round. When candidates are tied for lowest plurality score we break ties in lexicographically. The rule returns the lottery that assigns all probability to the single candidate that was never eliminated; $U(W(X))$ where $|W(X)|=1$.
Black’s rule is an example of a rule that is not a scoring rule, tournament rule, or weighted-tournament rule, but is still Condorcet-consistent.
Definition 15 (Black’s Rule).
If the profile $X$ admits a Condorcet winner $c$, then let $W(X)={c}$. Otherwise, if there is no Condorcet winner, let $W(X)$ be the subset of candidates with maximum Borda score $B(c)$. The probabilistic Black’s rule returns the lottery $U(W(X))$.
Appendix B Rule Preservation
Plurality and Borda are scoring rules, which are necessarily computable from a rank frequency embedding. However neither rule is preserved by the tournament embedding. Plurality is known not to be preserved by the weighted tournament either, but for Borda this remains an open question.
Plurality
Plurality is a scoring rule, and therefore necessarily computable from a rank frequency embedding. It requires only one column of information from the rank frequency matrix, representing how often each candidate is ranked first by a voter. By contrast, Plurality is not preserved by the weighted tournament embedding, and therefore not by the tournament embedding either.
Theorem 1.
The weighted tournament embedding does not preserve Plurality.
Proof.
$X_{1}=(a\succ b\succ c),(b\succ a\succ c),(c\succ a\succ b)$, $X_{2}=(a\succ b\succ c),(a\succ b\succ c),(c\succ b\succ a)$. ∎
Corollary 1.
The tournament embedding does not preserve Plurality.
Borda
Theorem 2.
The tournament embedding does not preserve Borda.
Proof.
$X_{1}=(a\succ b\succ c),(b\succ a\succ c),(b\succ c\succ a)$, $X_{2}=(a\succ b\succ c),(a\succ b\succ c),(a\succ b\succ c)$. ∎
Challenge.
Does the weighted tournament embedding preserve Borda?
Copeland
Copeland is the only probabilistic social choice function we consider that is preserved by the tournament embedding, and hence by the weighted tournament as well. However, Copeland is not preserved by the rank frequency embedding.
Theorem 3.
The rank frequency embedding does not preserve Copeland.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ a\succ d\succ c),(d\succ c\succ b\succ a)$. ∎
Schulze and Simpson-Kramer are weighted-tournament rules that are not preserved by the tournament or rank frequency embedding.
Schulze
Theorem 4.
The rank frequency embedding does not preserve Schulze.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ a\succ d\succ c),(d\succ c\succ b\succ a)$. ∎
Theorem 5.
The tournament embedding does not preserve Schulze.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$. ∎
Simpson-Kramer (Maximin)
Theorem 6.
The rank frequency embedding does not preserve Simpson-Kramer.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ a\succ d\succ c),(d\succ c\succ b\succ a)$. ∎
Theorem 7.
The tournament embedding does not preserve Simpson-Kramer.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ a\succ d),(d\succ c\succ a\succ b)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ c\succ a\succ d),(c\succ a\succ b\succ d)$. ∎
IRV
Theorem 8.
The rank frequency embedding does not preserve IRV.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ a\succ d\succ c),(d\succ c\succ b\succ a)$. ∎
Theorem 9.
The tournament embedding does not preserve IRV.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ c\succ b)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$. ∎
Challenge.
Does the weighted tournament embedding preserve IRV?
Black’s Rule
Theorem 10.
The rank frequency embedding does not preserve Black’s Rule.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ a\succ d\succ c),(d\succ c\succ b\succ a)$. ∎
Theorem 11.
The tournament embedding does not preserve Black’s Rule.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ a\succ d),(d\succ c\succ a\succ b)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ c\succ a\succ d),(c\succ a\succ b\succ d)$. ∎
Challenge.
Does the weighted tournament embedding preserve Black’s Rule?
Appendix C Training Loss for PSCFs
In this Appendix we give the graphs of validation loss for the other rules studied in this paper.
Appendix D Scaling to Fewer Voters
This table gives the results of our scaling experiments to a smaller number of voters.
$T_{RF}$ $T_{WT}$ $T_{T}$
Plurality 0.004252 0.148373 0.146005
Borda 0.004678 0.004018 0.075410
Copeland 0.070351 0.051713 0.000698
Schulze 0.080797 0.066160 0.045688
Simpson-Kramer 0.081472 0.061241 0.046492
IRV 0.095856 0.114207 0.088140
Black’s Rule 0.033607 0.033087 0.044673
Table 6: Testing generalization of our models trained to learn PSCFs on uniformly random profiles of 7 candidates and 29 voters on random profiles with $n=9$ voters. Average L1 losses over 1,000 random profiles.
Appendix E Participation-Adjustment
In this section we give the validation losses for learning participation by retraining our models for a given voting rule.