### v2

# DeepVoting: Learning and Fine-Tuning Voting Rules with Canonical Embeddings

Abstract
Aggregating agent preferences into a collective decision is an important step in many problems (e.g., hiring, elections, peer review) and across areas of computer science (e.g., reinforcement learning, recommender systems). As Social Choice Theory has shown, the problem of designing aggregation rules with specific sets of properties (axioms) can be difficult, or provably impossible in some cases. Instead of designing algorithms by hand, one can learn aggregation rules, particularly voting rules, from data. However, prior work in this area has required extremely large models or been limited by the choice of preference representation, i.e., embedding. We recast the problem of designing voting rules with desirable properties into one of learning probabilistic functions that output distributions over a set of candidates. Specifically, we use neural networks to learn probabilistic social choice functions. Using standard embeddings from the social choice literature we show that preference profile encoding has significant impact on the efficiency and ability of neural networks to learn rules, allowing us to learn rules faster and with smaller networks than previous work. Moreover, we show that our learned rules can be fine-tuned using axiomatic properties to create novel voting rules and make them resistant to specific types of “attack”. Namely, we fine-tune rules to resist a probabilistic version of the No Show Paradox.
1 Introduction
Computational Social Choice (COMSOC) and Algorithmic Game Theory (AGT) focus heavily on the design and analysis of mechanisms for collective decision making. Canonically, agents arrive with individual preferences over a set of alternatives or outcomes, and a mechanism aggregates these preferences into a shared choice (voting and selection) or allocation (matching and auctions) (Shoham and Leyton-Brown 2008). The goal is to design mechanisms with certain desirable properties, characterized by axioms; i.e. optimizing a particular objective or satisfying certain constraints.
A central result in Social Choice is Arrow’s General Impossibility Theorem (Arrow 1963), which identifies a set of axioms that no collective choice mechanism can satisfy simultaneously. Following Arrow, decades of research has produced myriad theorems showing which axioms are satisfied by which mechanisms and which lead to an impossibility results (Sen 2018), including optimality, computational complexity, and strategyproofness (Brandt et al. 2016).
Finding rules that satisfy a given set of axioms can be difficult, especially when it is unknown if such a rule exists. Hence, recent work has turned to machine learning techniques to design novel mechanisms. This idea has been applied to auctions, voting rules, matchings, and beyond (Xia 2013; Sandholm 2003; Curry et al. 2022; Ravindranath et al. 2021). Previous work on learning voting rules has been hampered by technical challenges, including extremely large/sophisticated neural nets (Anil and Bao 2021), limited data (Burka et al. 2022), or failure to account for the full consequences of the design choices (Firebanks-Quevedo 2020).
We improve the learning of existing and novel voting rules using common embeddings from the social choice literature. These embeddings enable faster learning with fewer parameters and scale to larger voter populations with better accuracy. Anil and Bao (2021) observed that using a multi-layer perceptron (MLP), i.e., a neural network, to learn voting rules was hampered by the network’s fixed input size, thus requiring more sophisticated architectures to permit scaling. Our embeddings reduce the input size of our neural net, greatly reducing the number of model parameters.
Sen (2018) observed that aggregation mechanisms can be described by what information they use or ignore from preference profiles. A key challenge in designing neural networks for voting rules is understanding how to handle different numbers of voters or candidates, and several embeddings proposed in the social choice literature may provide a solution. In particular, embeddings whose size is independent of the number of voters can improve learning, but the choice of embedding must correspond to the learning objective. Like any compression algorithm, embeddings can be lossy and impose a bound on the learnability of rules and axioms. A key contribution of our work is a more complete understanding of the relationship between the choice of embedding and the resulting learnability and efficiency of the mechanisms. Our experiments inspire new theoretical questions about the information preservation and choice of embeddings.
Often, when people think of voting they think of classical deterministic rules that take in preferences over a small set of candidates and return a single winner (Zwicker 2016; Taylor 2005). However, the full space of social choice mechanisms is much richer with mechanisms varying by the data types of their inputs and outputs; voters may give approval ballots, rankings, scores, or weightings to different candidates; the outcome of the mechanism may be a single winning candidate, collection of winners, or ordering of the candidates.
We study probabilistic social choice functions (PSCFs) which take a profile over candidates and return a lottery (probability distribution) over the candidates (Brandt 2017). Unlike single-winner rules, PSCFs provide a natural connection between the discrete nature of rules and axioms and the continuous loss functions for training based on divergences between distributions. We use the L1 loss, the sum of point-wise absolute differences (taxicab distance) between the source (neural network result) and target distribution (voting rule) (Abu-Mostafa, Magdon-Ismail, and Lin 2012).
We explore how well we can learn standard voting rules with common embeddings, testing against profiles both in and out of distribution. We then address the challenge of fine-tuning these networks to improve their axiomatic properties. We focus on the No Show Paradox in which a voter can induce an outcome they prefer by not voting (Moulin 1988). Single-winner Plurality, Borda, and Simpson-Kramer are known to satisfy this Participation axiom, though many other common voting rules are vulnerable to it (Zwicker 2016). We take models for PSCFs and fine-tune them using a loss function that adds in a continuous relaxation of the Participation axiom, showing rules can be refined to be more resistant to the paradox and maintain accuracy.
We choose the Participation axiom because it is an inter-profile axiom, which requires reasoning about counter-factuals on what the preference profile could have been had the voters behaved differently. Inter-profile axioms are particularly challenging for learning from data as we must consider many different profiles, e.g., all $m!$ manipulations, in training. It also requires that the model be able to take profiles of different sizes (differing by one voter) as input, which our embeddings enable us to do. Since abstention can be a strategic behavior by voters, our work is closely related to Automated Mechanism Design, which aims to create desirable mechanisms for strategic agents and “shifts the burden of design from man to machine” (Sandholm 2003).
Critiques of using machine learning methods for voting rules include (1) most voting rules are simple to compute, why complicate it? and (2) how do we explain these rules if they are the output of a network? In regards to (1) we take an engineering approach: the first part of this paper is a study on how effectively we can learn these rules, so that we can then judge how well our more participation-proof rules work. In answer to (1) and (2) we agree that for many cases a direct implementation of the rule may be better. However, in some cases like recommender systems (Aird et al. 2024; Patro et al. 2020), where we want to optimize an objective, and limit our downsides, one may be okay with using a less explainable rule. [^1] Note that run-time efficiency is a key metric for recommender systems, and inference of our models is extremely fast. Ultimately, we want to learn novel rules that sit at the empirical Pareto front of an optimization criteria (e.g., top-cycle or Condorcet consistency) and resistance to forms of attack (manipulation, strategic abstention). This work is a concrete step in that direction, showing the limits of learning, and pointing out ways forward.
Contribution.
We (1) explicitly characterize which common embeddings from Social Choice are able to retain desirable properties and can be used to learn popular voting rules; (2) demonstrate, for the first time, that standard embeddings from Social Choice dramatically reduce the complexity and increase the efficiency of learning voting rules; (3) use transfer learning (fine-tuning) to add an axiomatic property to a learned voting rule, thereby making existing rules more resistant to strategic manipulation; and (4) provide strong evidence that training on Impartial Culture preferences teaches rules to generalize to additional preference distributions.
2 Related Work
Xia (2013) and Procaccia et al. (2009) proposed incorporating voting axioms into a machine-learning framework as a means of evaluating learned social choice mechanisms. In the space of auction design and matching there has been work on using neural nets for better mechanisms (Dütting et al. 2019; Pavlov 2011; Malakhov and Vohra 2008) including learning new types of auction mechanisms (Curry et al. 2022) as well as complex preference structures (Peri et al. 2021). More recently, the work of Ravindranath et al. (2021) has looked at how to learn new allocation mechanisms that bridge the gap between stability (as compared to the deferred acceptance algorithm (Gale and Shapley 1962)) and strategyproofness (as compared to random serial dictatorship (RSD) (Aziz, Brandt, and Brill 2013)). While the work of Ravindranath et al. (2021), Firebanks-Quevedo (2020), and most recently Anil and Bao (2021), has shown promise for learning mechanisms, these efforts do not closely consider the role of embeddings. Armstrong (2025) considered the impact of embeddings on rule learnability across a wide range of preference distributions but used a very limited network size and provided little subsequent analysis.
While formal proposals to learn voting rules date back over a decade (Xia 2013), considerable attention to learning voting rules has increased in recent years. Kujawska, Slavkovik, and Rückmann (2020) and Burka et al. (2022) used several common machine learning methods to mimic existing voting rules. However, both of these works overlooked the importance of the choice of embedding in the role of learning, finding that certain rules were “easier” to learn but not theoretically characterizing why certain embeddings maintain properties, as we do. Subsequently, Anil and Bao (2021) showed that PIN architectures offer better generalization to larger numbers of voters. We build on this work by showing that we can achieve high accuracy efficiently with smaller MLPs by using specific embeddings.
Procaccia et al. (2009) showed that positional scoring rules are efficiently PAC learnable, but learning pairwise comparison-based voting rules requires an exponential number of samples. While we do not escape the asymptotic limits, we examine two embeddings based on tournament graphs that facilitate more efficient learning of pairwise-comparison based rules. Firebanks-Quevedo (2020) uses one measure of optimality (Condorcet consistency) and strategyproofness for learning. However, as we show, the chosen embedding in that work cannot learn strategyproofness, leading to poor results. Finally, Wilson (2019) focus on learning a voting rule given pair-wise relations and properties that must hold for the optimization criteria. However, they focused on the possibility of learning these functions and does not employ any ML techniques.
The loss function chosen by Armstrong and Larson (2019) was a function of the profile and outcome, and thus could learn a rule but not inter-profile axioms such as Participation. Recently, Mohsin et al. (2022) focused on the problem of designing and/or learning fair and private rules using random forests and a subset of embeddings we study, proving that under differential privacy there is an upper bound on the trade-off between group fairness and efficiency. Learning voting rules bears some similarity to the well studied area of learning to rank (L2R) from the machine learning literature (Cao et al. 2007). L2R is concerned with accurate recovery of the population preference and not the axioms or properties of the aggregation method itself (e.g., fairness). Indeed, one can think of our work enforcing inter-profile axioms on the learned aggregation procedures as an important step.
3 Preliminaries
Agents and Preference Profiles
Let $V$ be a set of $n$ voters and $C$ a set of $m$ candidates. Each voter $i\in V$ reports a strict order $x_{i}$ over all candidates in $C$ as their ballot. We denote that $i$ strictly prefers $a$ over $b$ by $a\succ_{i}b$ for $a,b\in C$. There are $m!$ possible ballots, or ways to strictly order (permute) the candidates in $C$. A list of $n$ ballots, one for each voter, constitutes a profile $X=(x_{i})_{i\in V}$. Voter $i$ ranks candidate $a$ at position $x_{i}(a)\in[m]$, using $k]=\{1,\ldots,k\}$. Let $\mathcal{X}$ be the set of all possible profiles.
Probabilistic Social Choice Functions
[…]
3.1 Embeddings
Traditional feed-forward neural networks require a fixed-size input for learning and inference, corresponding to the size of their input layer (Goodfellow et al. 2016). If we were to learn voting rules using neural networks that take the entire profile as input, then not only does the input layer need to be large $(m\times n)$, but it also prevents scaling up as the number of voters grows without resorting to more complex models (sequential or PINs) which are significantly larger, harder to train, and slower for inference (Anil and Bao 2021). Similarly, if the number of voters shrinks, then the profile would have to be padded carefully to preserve performance. To learn rules that are agnostic to the number of voters we need embeddings of a fixed-size that retain relevant information for profiles with any number of voters. Naturally, different embeddings preserve different information from the original profile, leading to different efficacy when learning different rules and axioms. Note that most rules and axioms in the literature are defined for any positive number of voters, so we would like our learned mechanisms to be similarly agnostic.
Figure 1: Each of the three embeddings derived from a ballot profile. Note that the size of the profile (a) grows in $O(mn)$, while each of our embeddings grows with $O(m^{2})$, which is far smaller when $m<<n$. However, our embeddings do not always preserve all of the information in the original profile.
An embedding $T$ is a function $T:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ mapping profiles to some codomain $\mathcal{X}^{\prime}$. The embeddings we are concerned with are many-to-one mappings. This means multiple different profiles may have the same embedding, i.e. T(X)=T(X~)T(X)=T($\tilde${X}) for some $X,\tilde{X}\in\mathcal{X}$ where $X\neq\tilde{X}$. In other words, $T$ will not be reversible, and $T(X)$ will not preserve all information about $X$. We denote by $\mathcal{F}^{\prime}$ the set of all probabilistic functions of the form $f^{\prime}:\mathcal{X}^{\prime}\rightarrow\Delta(C)$. Note that while we designate $\mathcal{X}$ to always contain strict orders over candidates, the structure of $\mathcal{X}^{\prime}$ will be different for different embeddings. The following three embeddings are drawn from the voting literature, but are not commonly recognized as embeddings in the machine learning literature.
Definition 1 (Tournament Embedding).
The tournament embedding $T_{T}$ yields a $m\times m$ matrix $M$ where $M[j,k]=1$ if a majority of voters prefer $j\succ_{i}k$, $M[j,k]=0$ if a majority prefer $k\succ_{i}j$, and $M[j,k]=\frac{1}{2}$ if an equal number of voters prefer each candidate (when $n$ is even), for candidate pairs $j,k\in C$.
Definition 2 (Weighted Tournament Embedding).
The weighted tournament embedding $T_{WT}$ yields a $m\times m$ matrix $M$ where $M[j,k]=|\{i\in V:j\succ_{i}k\}|$ for $j,k\in C$.
Observe that $T_{WT}$ contains strictly more information about the original profile than $T_{T}$ as the tournament can be computed from the weighted tournament.
Definition 3 (Rank Frequency Embedding).
The rank frequency embedding $T_{RF}$ yields a $m\times m$ matrix $M$ of how many voters rank each candidate $c\in C$ in each position $k\in[m]$ where $M[c,k]=|\{i\in V\text{ s.t. }\succ_{i}^{c}=k\}|$.
There is a tension in the literature between rules that use positional information, like scoring rules, and those that rely on majoritarian or pairwise comparison information, like tournament rules (Brandt, Brill, and Harrenstein 2014). Note $T_{RF}$ maintains positional information while $T_{WT}$ and $T_{T}$ are majoritarian. At times we also refer to a concatenation of all three embeddings which we refer to as $T_{CO}$.
Definition 4 (Combined Embedding).
The combined embedding $T_{CO}$ is the $3m^{2}$ concatenation of $T_{RF},T_{T},T_{WT}]$.
3.2 Probabilistic Social Choice Functions
We now define our PSCFs. Where necessary, we always break ties lexicographically. Definitions for Plurality, Schulze, Instant Runoff Voting (IRV) and Black’s rule can be found in the Appendix. Two rules that are typically classified as scoring rules, Borda and Plurality. The outcome of any scoring rule can be exactly computed from $T_{RF}$.
Definition 5 (Borda).
The Borda score of candidate $c\in C$ from profile $X$ is $B(c)=\sum\limits_{i\in V}(m-x_{i}(c))$. Let $W(X)=\operatorname*{arg\,max}\limits_{c\in C}B(c)$ be the subset of candidates with maximum Borda score. The probabilistic Borda rule returns the lottery $U(W(X))$ (Referred to as Borda Max by Endriss (2017)).
The rest of our rules are not scoring rules. Copeland is a tournament rule since it’s outcome can be computed directly from $T_{T}$ (Brandt, Brill, and Harrenstein 2014).
Definition 6 (Copeland).
The Copeland score of candidate $c\in C$ from profile $X$ is the number of other candidates it beats in pairwise competition plus $\frac{1}{2}$ times the number of other candidates it ties with in direct competition (if $n$ is even). Let $W(X)$ be the subset of candidates with maximum Copeland score on profile $X$. The probabilistic Copeland rule returns the lottery $U(W(X))$.
The Simpson-Kramer (Maximin) and Schulze rules are each computed from $T_{WT}$. We will call these weighted-tournament rules. Let $G_{X}(C,E)$ be the directed tournament graph with edges corresponding to all positive values of the tournament matrix induced by $X$. Let each directed edge $(a,b)\in E$ have weight $d(a,b)=|i\in V:a\succ_{i}b|$.
Definition 7 (Simpson-Kramer).
Let $W$ be the subset of candidates whose maximum weight incoming edge is minimal in $G_{X}$. The probabilistic Simpson-Kramer rule returns the lottery $U(W(X))$.
Some, but not all, of the rules listed above are Condorcet-consistent, meaning that they place all probability mass on the Condorcet winner whenever one exists. A Condorcet winner is a candidate who beats all other candidates in pairwise competition, which can be inferred from $T_{T}$ or $T_{WT}$.
4 PSCF Preservation Under Embedding
We are concerned with what information is preserved by embeddings, and whether this information is sufficient to implement PSCFs, i.e. to learn them perfectly.
Definition 8 (PSCF Preservation).
A PSCF $f:\mathcal{X}\rightarrow\Delta(C)$ is preserved by embedding $T:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ if $\exists f^{\prime}:\mathcal{X}^{\prime}\rightarrow\Delta(C)$ such that $f^{\prime}(T(X))=f(X)$ for all profiles $X\in\mathcal{X}$.
Proposition 1 says that for an embedding $T$ to preserve a PSCF, there cannot be two profiles with the same embedding under $T$ for which the PSCF returns different lotteries.
Proposition 1.
Embedding $T$ preserves PSCF $f$ if and only if $f(X^)T(X)=T(\hat{X})\Rightarrow$ $f(\hat{X})$ for all $X,\hat{X}\in\mathcal{X}$.
Some embeddings preserve strictly more information than others. For instance, $T_{WT}$ preserves all information necessary to compute $T_{T}$ from a profile. This implies that if $T_{T}$ preserves a function $f$, then $T_{WT}$ must preserve $f$ as well.
Proposition 2.
Suppose that for $T:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ there exist $T_{1}:\mathcal{X}\rightarrow\hat{\mathcal{X}}$ and $T_{2}:\hat{\mathcal{X}}\rightarrow\mathcal{X}^{\prime}$ such that $T(X)=T_{2}(T_{1}(X))$ for all $X\in\mathcal{X}$. Then for all $f\in\mathcal{F}$, $T$ preserves $f$ only if $T_{1}$ preserves $f$.
As we can compute $T_{T}$ from $_{WT}, TT$$_{T}$ can only preserve a PSCF if $T_{WT}$ does as well, the reverse does not hold. $T_{WT}$ may preserve PSCFs that are not preserved by $T_{T}$. If an embedding preserves a PSCF, then the PSCF is perfectly learnable from the embedding. Table 1 (green highlights) shows which PSCFs are preserved by these embeddings. See Appendix B for proofs of the negative results where PSCFs are not preserved. Each of these proofs consists of a counterexample with two profiles whose outcomes differ under the PSCF yet have the same embedding. Finally, many of our rules fall into Fishburn’s categorization: the winner of C1 rules (Copeland) can be computed using the information encoded in $T_{T}$, while the winner of C2 rules (Borda, Schulze, Simpson-Kramer, Black’s) can be computed from $T_{WT}$. A separate categorization of voting rules, positional scoring rules (Plurality, Borda), can be computed using only $T_{RF}$  (Brandt et al. 2016). IRV has recently been shown not to belong to either C1 or C2, and is not a positional scoring rule (Halpern, Hossain, and Tucker-Foltz 2024).
Figure 2: Validation Loss per epoch for each rule and embedding pair.
5 Learning Lotteries from PSCFs
First, we show that with suitable embedding we can learn PSCFs that generalize common voting rules using network architectures with few parameters. We train rule-embedding pairs separately for 32 combination of rule and embedding (3 embeddings and their concatenation) to compare their performance, and explore the rule-embedding tradeoff. Mohsin et al. (2022) use some of the same embeddings with XGBoost. However, we are the first to use them with MLPs, and hence they must be validated. MLPs also allow us to fine-tune these rules later in Section Section 6, which is not possible with XGBoost.
Experimental Setup
We train our PSCFs on profiles with $n=44$ voters and $m=11$ candidates. For all experiments, profiles are sampled from the impartial culture distribution – i.e., rankings are generated uniformly at random (Black 1958). Like Firebanks-Quevedo (2020), we use the Whalrus package to implement our voting rules. [^2] https://pypi.org/project/whalrus/
Embeddings afford three key advantages: (1) They reduce the size of the input layer of our network, which is fully connected, and therefore greatly reduce the number of model weights. All three embeddings compress the $n\times m$ profile to an $m\times m$ matrix representation, so the same MLP architecture can be used for all training runs. (2) When a rule is paired with an appropriate embedding, the embedding preserves all the information necessary to learn the rule and removes unnecessary information. (1) and (2) mean that we learn PSCFs faster and more accurately than previous work. (3) Input size no longer depends on the number of voters, which lends itself to better scaling. For $T_{RF}$ and $T_{WT}$ we normalize by dividing all elements in the embedding by $n$, e.g., the elements of $T_{WT}$ represent the fraction of voters who prefer one candidate to another $\frac{d(a,b)}{n}$ for $a,b\in C$.
We emulate the MLP architecture of Anil and Bao (2021), with 5 fully-connected layers, the first hidden layer with 200 nodes, then four with 120 nodes, TanH activation functions, and a Softmax layer for the output. [^3] Discussion of other setups is in Appendix C. The key difference is that our network takes in embedded profiles so the size of our input layer is $m^{2}$ compared to their $nm^{2}$. This brings our total number of model parameters down to $\approx$100K vs. millions. We train our models on a set of 100,000 randomly sampled profiles in batches of size 32 for 1000 epochs, for a total of 1.5M gradient steps. The use of embeddings also allows us to test our MLP model on larger voter profiles without increasing the size of the network, which Anil and Bao (2021) were unable to do for their MLP model. We trained each model on NVIDIA A100 GPUs using PyTorch, with each run taking $\approx$4 hours. We used the Adam optimizer for each run with an initial learning rate of 0.001, tuning on plateau (patience = 50, factor = 0.5, min_lr = $1e^{-5}$).
We refer to the L1 distance between our model output and the PSCF lottery on a profile as the rule loss. Rule losses presented in Table 1 are from a test set of 10,000 random profiles sampled independently of the training data. All models are trained to minimize rule loss for their PSCFs.
Learned PSCFs
Target Rule $_{RF} TT$$_{T}$ $T_{WT}$ $T_{CO}$
Plurality 0.0 0.129 0.136 0.001
Borda 0.165 0.048 0.018 0.042
Copeland 0.088 0.0 0.163 0.0
Schulze 0.071 0.043 0.067 0.042
Simpson-Kramer 0.071 0.043 0.062 0.043
IRV 0.159 0.105 0.159 0.161
Black’s Rule 0.046 0.044 0.165 0.037
Ranked Pairs 0.161 0.051 0.161 0.07
Table 1: L1 Loss for each embedding on test data sampled from the Impartial Culture for models targeting each rule using $(m=11,n=44)$. Shaded cells indicate the embedding contains sufficient information to learn the rule perfectly.
Table 1 gives final validation set results and Figure 2 shows plots of our validation losses during training, using 10,000 samples from a held-out validation set, to demonstrate the effectiveness of learning for each rule-embedding pair. Plurality learns rapidly using $T_{RF}$ and $T_{CO}$ with rule losses converging to zero quickly. Our other positional scoring rule, Borda, has significant trouble learning from $T_{RF}$ despite the embedding having enough information to compute the Borda winner. For Plurality, we see some learning from $T_{WT}$ and $T_{T}$, but learning quickly plateaus as the embeddings do not preserve all information needed to learn the rule, and so there is a non-zero lower bound to the error rate. The Copeland rule learns rapidly with the $T_{T}$ and $T_{CO}$, converging to near zero loss quickly since it is preserved. While any rule that can be computed from $T_{T}$ can also be computed from $T_{WT}$, what we see is that the Copeland rule loss falls far more slowly with $T_{WT}$, failing to reach the same loss as $T_{T}$ in our experiments after 1000 epochs. We make a similar observation for the Schulze rule. However, unlike Copeland, Schulze can be computed exactly from $T_{WT}$ but not $T_{T}$. This is why we see the loss with $T_{T}$ plateau at a nonzero value for Schulze. Looking at Table 1 we can see the variation of final loss across all embeddings. While $T_{CO}$ contains all the information, in some cases we are able to more effectively learn from smaller embeddings. This highlights the challenges of working with neural networks, the benefits of choosing the right embedding for the rule, and that embeddings containing more information do not always help learning.
5.1 Comparison to Single-Winner
We now test our PSCFs for their accuracy in identifying the unique winners of each rule (when they exist) to directly compare with the four rules (Plurality, Borda, Copeland, and Simpson-Kramer) of Anil and Bao (2021). For each rule, we sample profiles with unique winners, obviating the problems of tie-breaking, and select the candidate with the highest probability mass as the winner. [^4] This rejection sampling method of Anil and Bao (2021) eliminates profiles with multiple winners, which may introduce artifacts into the accuracy measures. Our PSCFs do not share this problem. Note that our MLP architecture is the same as Anil and Bao (2021), only differing with better embeddings, and results are given in Table 2. We can see that our models learn Plurality, Borda, Copeland, and Simpson-Kramer extremely well. Our performance for these four rules is on par (Plurality, Borda) or better (Copeland, Simpson-Kramer) with the results of Anil and Bao (2021), and in some cases we are able to outperform even their more complex PIN architectures. For example, we learn Copeland perfectly (1.0) whereas across their four architectures their best performing model is 0.83; Simpson-Kramer our models (0.913) strictly outperform all of theirs (best of 0.80). Hence by leveraging embeddings we are able to learn rules as well as or better with models orders of magnitude smaller.
Target Rule $_{RF} TT$$_{T}$ $T_{WT}$ $T_{CO}$
Plurality 0.999 0.355 0.289 0.997
Borda 0.086 0.826 0.934 0.844
Copeland 0.552 1.0 0.087 1.0
Schulze 0.719 0.903 0.735 0.914
Simpson-Kramer 0.72 0.902 0.79 0.913
IRV 0.118 0.446 0.118 0.106
Black’s Rule 0.767 0.866 0.088 0.88
Ranked Pairs 0.102 0.814 0.102 0.67
Table 2: Accuracy of our models on each embedding using test data sampled from the Impartial Culture using $(m=11,n=44)$. Most accurate embeddings are shown in bold.
5.2 Beyond Impartial Culture
We now consider the generality of our training distribution against other preference distributions. As the Impartial Culture (IC) provides orders where candidates are ranked uniformly at random, all possible profiles have a non-zero probability of occurring. That is, given sufficient data, IC will generate profiles that could have been generated by all other distributions. However, recent (Boehmer et al. 2024, 2022; Szufa et al. 2025) and older (Mattei and Walsh 2013, 2017) work in the COMSOC community has illustrated the need to test on a wide variety of synthetic and real world preference distributions to ensure generalization. We test the empirical merit of this fact by evaluating our networks trained on IC preferences on test sets sampled from a wide range of distributions, including real-world preference data, complete results and definitions of distributions are in Appendix E.
Target Rule IC IAC Urn Mall. SP PrefL
Plurality .001 .001 .018 .012 .031 .014
Borda .028 .03 .077 .093 .135 .109
Copeland .0 .0 .001 .001 .0 .002
Schulze .017 .016 .029 .031 .018 .042
SK .017 .016 .028 .027 .02 .039
IRV .163 .162 .163 .164 .182 .162
Black’s .022 .023 .066 .085 .018 .103
RP .06 .06 .061 .06 .001 .063
Table 3: L1 Loss across preference distributions of networks trained using Impartial Culture and the $T_{CO}$ embedding.
Table 3 shows the loss of networks trained on $T_{CO}$ tested on 10,000 profiles from each other distribution. We include this evaluation for networks trained on other embeddings in Appendix E. In all cases, rules trained on $T_{CO}$ are able to generalize very effectively to new distributions. On highly structured preferences, such as the Single-Peaked distribution, some rules (e.g., Black’s, Ranked Pairs) have lower loss than on Impartial Culture preferences.
6 Resisting the No Show Paradox
Figure 3: Validation losses for models trained to learn PSCFs using $T_{CO}$ with $m=11,n=44$ and retrained to learn Participation. Combined loss shows the sum of rule and Participation Loss. Most rules converge quickly.
PSCFs based on voting rules can be vulnerable to the No Show Paradox, where a voter prefers the outcome yielded by a rule when they do not vote, giving an incentive to abstain. A rule for which this cannot occur is said to satisfy the Participation axiom. We now employ transfer learning, taking our trained models from Section 5 and retraining them with a loss function that adds a term for Participation loss.
For all definitions below, let $P_{X}$ be a probability distribution derived from profile $X$ by some PSCF $f$ (implicit), and let $P_{X}(c)$ be the probability assigned to candidate $c$. Where the specific profile is not relevant, we will denote simply by $P(c)$ the probability assigned to candidate $c$ by a lottery $P$. We use stochastic dominance to model a voter’s preference between two lotteries based on their preference order over candidates to define Participation for PSCFs.
Definition 9 (Stochastic Dominance).
Let $\sigma$ be an ordering (or permutation) over the set of candidates $C$, and let $\sigma[k]$ be the $k^{th}$ element of $\sigma$ for $k\in[m]$. Given two lotteries $P$ and $Q$ over $C$, $P$ stochastically dominates $Q$ with respect to $\sigma$ if for all $k\in[m]$, $\sum\limits_{l\leq k}P(\sigma[l])\geq\sum\limits_{l\leq k}Q(\sigma[l])$.
We say that a voter’s abstention leads to an outcome ($P$) they prefer if the new outcome stochastically dominates the outcome ($Q$) that would derive from the true profile, with respect to the voter’s ordering of the candidates $\sigma=x_{i}$. We want our PSCF immune to strategic abstentions.
Definition 10 (Participation).
A PSCF $f$ obeys Participation if, for all profiles, every voter prefers the outcome under $f$ when they vote their true preference to the outcome under $f$ when they abstain (i.e. removed). We say that a voter prefers the outcome $Q$ from voting truthfully over the lottery $P$ from abstaining if $Q$ stochastically dominates $P$.
Since Participation is a binary condition for a PSCF, to learn PSCFs that resist the No Show Paradox, we define a non-binary loss function based on stochastic dominance.
Definition 11 (Stochastic Dominance Loss).
Given ordering $\sigma$ over $C$, a lottery $P$, and a reference lottery $Q$, we say that the stochastic dominance loss is zero if $P$ stochastically dominates $Q$. If $P$ does not stochastically dominate the reference lottery $Q$, then the loss is equal to $L(P|\sigma,Q)=\max\limits_{k\in[m]}(\sum\limits_{l\leq k}Q(\sigma[l])-\sum\limits_{l\leq k}P(\sigma[l]))$, i.e. the largest difference between the sums of prefixes of the lotteries over all prefixes when the distributions’ supports are ordered by $\sigma$.
Definition 12 (Participation Loss).
Given a profile $X$, Let $P^{i}_{X}$ be the lottery under $f$ when voter $i$ abstains and all others vote truthfully, and let $Q_{X}$ be the lottery under $f$ when voting truthfully. $L(f,X)=\max\limits_{i\in V}L(P^{i}_{X}|\sigma,Q_{X})$
Experimental Setup
Fine-tuning uses the same architectures and setup as the initial training. However, we retrain on 1056 random profiles with 11 candidates and 44 voters as, like most manipulations, the No Show Paradox is more likely to occur with fewer voters (Xia and Conitzer 2008). Changing the numbers of voters, without padding the profile, is a benefit of our embeddings. We add Participation Loss and the original rule loss for each profile and retrain for 40 total epochs taking about 12 hours each. Using fewer voters for training is also more computationally efficient, which is important as computing losses based on $n$ alternative profiles for each profile increases the runtime by $O(n)$. This is a major challenge for all inter-profile axioms that involve counterfactual comparisons as it determines how many different profiles must be considered determine if an axiom is satisfied (Schmidtlein 2022; Schmidtlein and Endriss 2023).
Before FT After FT
Target Rule Rule Part. Rule Part.
Plurality 0.0005 0.0713 0.0294 0.0000
Borda 0.0381 0.0000 0.0432 0.0000
Copeland 0.0000 0.2220 0.0051 0.0000
Schulze 0.0304 0.2152 0.0265 0.0000
Simpson-Kramer 0.0330 0.0015 0.0323 0.0000
IRV 0.1629 0.0000 0.1629 0.0000
Black’s Rule 0.0309 0.0000 0.0488 0.0000
Ranked Pairs 0.0699 0.0000 0.0797 0.0000
Table 4: Loss of our $T_{CO}$ models before and after fine-tuning with Participation Loss using $(m=11,n=44)$.
Participation-Adjusted PSCFs
Figure 3 gives the training loss per epoch while Table 4 gives both the Rule Loss and Participation Loss of rules before and after fine-tuning evaluated on a disjoint 1056 profiles from the training set; additional results for all embeddings are in Appendix F. These results are interesting in several ways. First, the single-winner versions of Borda, Plurality, and Simpson-Kramer resist the No Show Paradox, but only our learned Borda PSCF is resistant, with others showing some loss before fine-tuning. The rest of our rules are known to suffer from the paradox (Pérez 2001), and it is known that Condorcet-consistency is incompatible with Participation when there are at least 4 candidates and 12 voters (Brandt, Geist, and Peters 2017), although the paradox does not arise frequently, only in about 4% of profiles (Brandt, Hofbauer, and Strobel 2019). This leads to one of the most interesting results, we see both Copeland and Schulze, Condorcet Consistent rules, able to be fine-tuned in a way that mostly preserves the rule loss, but is also (empirically) immune to the no show paradox. While these are only small scale test, they point an intriguing way forward for future research.
7 Conclusions and Future Work
We have shown that not only can we efficiently and accurately learn known PSCFs from preference data, but also that we can fine-tune these rules in order to improve them in ways that, to date, have not be possible through traditional algorithmic design methods. We have highlighted the importance of the choice of embedding on the efficiency of learning and quality of the learned rules. For Participation, we saw that our models trained only on rules did reasonably at satisfying the axiom. After fine tuning, all rules are empirically Participation-proof with minimal loss in rule performance. These adjustment for improving axiomatic properties would be much more difficult, or intractable, to design by hand. Interestingly, we showed none of the tested embeddings retain all the information necessary for IRV, and we see that IRV is indeed the hardest rule to learn across all our testing, reinforcing the importance of embedding/target selection. It remains to be seen whether other embeddings can be designed, of size $m\times m$ or smaller, that outperform the embeddings we took from the social choice literature. Different embeddings may be beneficial in particular for rules whose outcomes are NP-Hard to compute.
Acknowledgments
Matone, Armstrong, and Mattei were supported in part by NSF Awards IIS-RI-2007955, IIS-III-2107505, IIS-RI-2134857, IIS-RI-2339880 and CNS-SCC-2427237 as well as the Harold L. and Heather E. Jurist Center of Excellence for Artificial Intelligence at Tulane University and the Tulane University Center of Excellence for Community-Engaged Artificial Intelligence (CEAI). Ben Abramowitz was supported by the NSF under Grant #2127309 to the Computing Research Association for the CIFellows Project.
Appendix For: DeepVoting: Learning and Fine Tuning Voting Rules with Canonical Embeddings
Appendix A Additional Voting Rules
In this section we give full definitions of other voting rules we study.
Definition 13 (Plurality).
The Plurality score of candidate $c\in C$ from profile $X$ is $L(c)=|\{i\in V:x_{i}(c)=1\}|$. Let $W(X)=\operatorname*{arg\,max}\limits_{c\in C}L(c)$ be the subset of candidates with maximum Plurality score. The probabilistic Plurality rule returns the lottery $U(W(X))$.
Definition 14 (Schulze).
For each path from $a$ to $b$ in $G_{X}$, we let the strength of the path be the minimum weight edge in that path. For each pair of candidates $a,b\in C$ with a path from $a$ to $b$, we let $p(a,b)$ be the maximum strength of any path from $a$ to $b$, and let $p(a,b)=0$ otherwise. Finally, let $W(X)=\{a\in C:p(a,b)\geq p(b,a)\text{ for all }b\in C\}$. The probabilistic Schulze rule returns the lottery $U(W(X))$.
Instant Runoff Voting is not a scoring rule, but is defined by iteratively using plurality scores.
Definition 15 (Instant Runoff Voting (IRV)).
IRV is a deterministic, iterative voting rule that, in each of m1m-1 rounds, eliminates the candidate with the lowest plurality score and removes them from the preference orders of all voters before the next round. When candidates are tied for lowest plurality score we break ties in lexicographically. The rule returns the lottery that assigns all probability to the single candidate that was never eliminated; $U(W(X))$ where $|W(X)|=1$.
Black’s rule is an example of a rule that is not a scoring rule, tournament rule, or weighted-tournament rule, but is still Condorcet-consistent.
Definition 16 (Black’s Rule).
If the profile $X$ admits a Condorcet winner $c$, then let W(X)=cW(X)={c}. Otherwise, if there is no Condorcet winner, let $W(X)$ be the subset of candidates with maximum Borda score $B(c)$. The probabilistic Black’s rule returns the lottery $U(W(X))$.
Appendix B Rule Preservation
Plurality and Borda are scoring rules, which are necessarily computable from a rank frequency embedding. However neither rule is preserved by the tournament embedding. Plurality is known not to be preserved by the weighted tournament either.
$T_{RF}$ $_{WT} TT$$_{T}$
Plurality $\times$ $\times$
Borda $\times$
Copeland $\times$
Schulze $\times$ $\times$
Simpson-Kramer $\times$ $\times$
IRV $\times$ $\times$ $\times$
Black’s Rule $\times$ $\times$
Table 5: PSFC preservation under embedding.
Plurality
Plurality is a scoring rule, and therefore necessarily computable from a rank frequency embedding. It requires only one column of information from the rank frequency matrix, representing how often each candidate is ranked first by a voter. By contrast, Plurality is not preserved by the weighted tournament embedding, and therefore not by the tournament embedding either.
Theorem 1.
The weighted tournament embedding does not preserve Plurality.
Proof.
$X_{1}=(a\succ b\succ c),(b\succ a\succ c),(c\succ a\succ b)$, $X_{2}=(a\succ b\succ c),(a\succ b\succ c),(c\succ b\succ a)$. ∎
Corollary 1.
The tournament embedding does not preserve Plurality.
Borda
Theorem 2.
The tournament embedding does not preserve Borda.
Proof.
$X_{1}=(a\succ b\succ c),(b\succ a\succ c),(b\succ c\succ a)$, $X_{2}=(a\succ b\succ c),(a\succ b\succ c),(a\succ b\succ c)$. ∎
Copeland
Copeland is the only probabilistic social choice function we consider that is preserved by the tournament embedding, and hence by the weighted tournament as well. However, Copeland is not preserved by the rank frequency embedding.
Theorem 3.
The rank frequency embedding does not preserve Copeland.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ a\succ d\succ c),(d\succ c\succ b\succ a)$. ∎
Schulze and Simpson-Kramer are weighted-tournament rules that are not preserved by the tournament or rank frequency embedding.
Schulze
Theorem 4.
The rank frequency embedding does not preserve Schulze.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ a\succ d\succ c),(d\succ c\succ b\succ a)$. ∎
Theorem 5.
The tournament embedding does not preserve Schulze.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$. ∎
Simpson-Kramer (Maximin)
Theorem 6.
The rank frequency embedding does not preserve Simpson-Kramer.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ a\succ d\succ c),(d\succ c\succ b\succ a)$. ∎
Theorem 7.
The tournament embedding does not preserve Simpson-Kramer.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ a\succ d),(d\succ c\succ a\succ b)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ c\succ a\succ d),(c\succ a\succ b\succ d)$. ∎
IRV
Theorem 8.
The rank frequency embedding does not preserve IRV.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ a\succ d\succ c),(d\succ c\succ b\succ a)$. ∎
Theorem 9.
The tournament embedding does not preserve IRV.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ c\succ b)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$. ∎
Theorem 10.
The weighted tournament embedding does not preserve IRV.
This follows as a corollary to Halpern, Hossain, and Tucker-Foltz (2024). Thanks to Daniel Halpern for pointing this out!
Black’s Rule
Theorem 11.
The rank frequency embedding does not preserve Black’s Rule.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ d\succ a),(d\succ a\succ b\succ c)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ a\succ d\succ c),(d\succ c\succ b\succ a)$. ∎
Theorem 12.
The tournament embedding does not preserve Black’s Rule.
Proof.
$X_{1}=(a\succ b\succ c\succ d),(b\succ c\succ a\succ d),(d\succ c\succ a\succ b)$, $X_{2}=(a\succ b\succ c\succ d),(b\succ c\succ a\succ d),(c\succ a\succ b\succ d)$. ∎
Challenge 1.
Does the weighted tournament embedding preserve Black’s Rule?
Appendix C Other Tested Neural Network Setups
We conducted several experiments to explore the efficiency of learning and target larger profile dimensionalities without making significant changes to our models or embeddings. Experimenting with our model’s representational threshold, we found that with higher dimensionalities from our initial baseline of 7 candidates and 29 voters we were able to achieve faster and more consistent divergence with L1Loss and TanH activation.
With 15 candidates and 44 voters, we observe training loss decrease much more consistently than our ReLU baseline with lower variance and better generalization. Similarly, we find that utilization of a distribution-based loss like KLDiv or JensenShannon loss is inefficient when targeting a sparse output vector, and that even with 7 candidates these approaches see a severe reduction in the generalizability of our models.
While we did want to focus on smaller model architectures to more directly rely on our embeddings, we did experiment with more complex models. We tested two alternate networks which we observed to have below-satisfactory performance when compared to our base feed-forward network. The first was a simple network with a depth of 6 and a constant layer width of $n^{2}\times 3$, and the second had the same depth, but a funnel architecture which decomposed the width of each hidden layer from the input to the output layer to minimize information dropoff. Both networks underperformed when compared to our baseline.
Appendix D Scaling Number of Voters
Scaling With Number of Voters
Our embeddings give us the ability to work with profiles with different numbers of voters. Although we trained our models on profiles with 44 voters, we can test with larger and smaller numbers of voters to check generalizability. We test first on profiles with 199 votes (Table 6), and then again with only 13 voters (Table 7).
We see that there is a very mild increase in the loss across most rules and embeddings, though losses remain very similar to their value on profiles with 44 voters (see Table 1). In some cases there is no increase in loss: With the concatenated embedding, $T_{CO}$, both Copeland (with 13 voters) and Black’s rule (199 voters) have the same loss, as does. Loss even decreases when evaluating Black’s rule on profiles with 13 voters and $T_{T}$. As well, on all but 4 rule-embedding pairs, loss is lower using profiles with 13 voters than those with 199 voters. These results highlight the importance of choosing embeddings that fit the rule, corresponding to the learning objective. As a result of our embeddings our learned rules are able to generalize extremely effectively to profiles of varying sizes without the need for additional training.
Target Rule $_{RF} TT$$_{T}$ $T_{WT}$ $T_{CO}$
Plurality 0.04 0.133 0.142 0.058
Borda 0.165 0.059 0.039 0.057
Copeland 0.095 0.0 0.159 0.001
Schulze 0.08 0.057 0.082 0.057
Simpson-Kramer 0.083 0.057 0.078 0.057
IRV 0.166 0.112 0.166 0.165
Black’s Rule 0.061 0.054 0.166 0.045
Ranked Pairs 0.164 0.059 0.164 0.078
Table 6: L1 Loss on 512 Impartial Culture profiles for networks trained on each embedding 11 candidates and 199 voters.
Target Rule $_{RF} TT$$_{T}$ $T_{WT}$ $T_{CO}$
Plurality 0.023 0.119 0.131 0.019
Borda 0.166 0.051 0.024 0.047
Copeland 0.095 0.001 0.159 0.0
Schulze 0.083 0.05 0.082 0.049
Simpson-Kramer 0.079 0.049 0.076 0.05
IRV 0.158 0.108 0.158 0.16
Black’s Rule 0.055 0.042 0.166 0.037
Ranked Pairs 0.165 0.055 0.165 0.072
Table 7: L1 Loss on 512 Impartial Culture profiles for networks trained on each embedding 11 candidates and 13 voters.
Appendix E Testing Novel Distributions
This section contains a brief definition of each distribution we tested our networks on and additional results showing the result of testing our networks on each embedding. We include one table for each embedding, showing the effectiveness of the learned embedding-rule pair on data sampled from each test distribution.
Preference Distributions
We train all of our results only on the Impartial Culture which generates profiles with candidates ranked uniformly at random, however we test our networks on each of the following:
  • Impartial Culture (IC) Each unique preference order is equally likely, regardless of which orders any other voters have selected  (Guilbaud 1952).
  • Impartial Anonymous Culture (IAC) Preferences profiles are generated collectively, rather than as individual preference orders. Each multi-set of preference orders (i.e. a preference profile) is equally likely to be generated, making voter identities irrelevant  (Gehrlein and Fishburn 1976).
  • Mallows Preference orders are noisy estimates of some reference ranking $r$, with the amount of noise related to a parameter $\phi$  (Mallows 1957). A value of $\phi=0$ results in all voters have identical preferences while $\phi=1$ results in the Impartial Culture distribution. For each profile we sample $\phi\in[0,1]$ uniformly at random using the Mallow’s distribution described by  Boehmer et al. (2021).
  • Urn All $m!$ preference orders exist in an “urn.” Each voter decides their ranking by sampling a ranking from the urn. Once a ranking is selected, $\alpha!$ copies of it are added to the urn  (Eggenberger and Pólya 1923). For each profile we sample $\alpha$ from a Gamma distribution with shape parameter $k=0.8$ and scale parameter $\theta=1$ as described by Boehmer et al. (2021).
  • Single-Peaked There is some global ordering of alternatives. Each voter has some favourite alternative and prefers all alternatives closer to their favourite over those further away. We sample single-peaked profiles from Walsh’s distribution (Walsh 2015).
  • PrefLib An online repository containing profiles corresponding to real human preferences expressed across many domains  (Mattei and Walsh 2013). We use all complete profiles with strict orders and $m\geq 11$ candidates. For profiles with greater than $11$ candidates we form a profile by selecting a subset of candidates uniformly at random. This results in a test set of 6945 profiles.
Loss and Accuracy
The following tables show L1 Loss and test accuracy on 10,000 test profiles (except in the case of PrefLib) sampled from each of the above distributions. We exclude from our data profiles where any rule results in multiple, tied winners.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.0 0.0 0.032 0.002 0.05 0.021
Borda 0.165 0.165 0.162 0.156 0.179 0.158
Copeland 0.088 0.088 0.109 0.116 0.181 0.112
Schulze 0.071 0.071 0.094 0.106 0.018 0.101
SK 0.071 0.071 0.094 0.106 0.176 0.104
IRV 0.159 0.16 0.163 0.165 0.182 0.165
Black’s Rule 0.046 0.046 0.076 0.079 0.171 0.075
Ranked Pairs 0.161 0.162 0.164 0.165 0.182 0.166
Table 8: L1 Loss across preference distributions of networks trained using Impartial Culture preferences and the $T_{RF}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.999 1.0 0.851 0.902 0.74 0.883
Borda 0.086 0.085 0.111 0.116 0.0 0.136
Copeland 0.552 0.554 0.408 0.431 0.0 0.379
Schulze 0.719 0.716 0.507 0.498 0.931 0.444
SK 0.72 0.716 0.502 0.489 0.017 0.431
IRV 0.118 0.114 0.097 0.092 0.0 0.092
Black’s Rule 0.767 0.765 0.574 0.596 0.048 0.588
Ranked Pairs 0.102 0.097 0.097 0.091 0.0 0.086
Table 9: Test Accuracy across preference distributions of networks trained using Impartial Culture preferences and the $T_{RF}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.129 0.131 0.084 0.054 0.09 0.065
Borda 0.048 0.05 0.06 0.007 0.033 0.019
Copeland 0.0 0.0 0.0 0.0 0.0 0.0
Schulze 0.043 0.042 0.025 0.005 0.001 0.01
SK 0.043 0.043 0.026 0.005 0.006 0.01
IRV 0.105 0.105 0.099 0.082 0.086 0.084
Black’s Rule 0.044 0.045 0.028 0.005 0.008 0.011
Ranked Pairs 0.051 0.051 0.036 0.017 0.001 0.02
Table 10: L1 Loss across preference distributions of networks trained using Impartial Culture preferences and the $T_{T}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.355 0.35 0.599 0.633 0.493 0.642
Borda 0.826 0.82 0.704 0.902 0.856 0.895
Copeland 1.0 0.999 1.0 1.0 1.0 0.981
Schulze 0.903 0.908 0.93 0.967 1.0 0.934
SK 0.902 0.906 0.931 0.975 0.999 0.933
IRV 0.446 0.448 0.489 0.531 0.555 0.537
Black’s Rule 0.866 0.86 0.914 0.96 0.994 0.94
Ranked Pairs 0.814 0.818 0.863 0.897 1.0 0.888
Table 11: Test accuracy across preference distributions of networks trained using Impartial Culture preferences and the $T_{T}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.136 0.137 0.122 0.109 0.094 0.113
Borda 0.018 0.019 0.048 0.078 0.018 0.071
Copeland 0.163 0.162 0.157 0.151 0.182 0.152
Schulze 0.067 0.067 0.118 0.138 0.137 0.133
SK 0.062 0.061 0.114 0.136 0.17 0.13
IRV 0.159 0.16 0.163 0.165 0.182 0.165
Black’s Rule 0.165 0.164 0.166 0.166 0.182 0.166
Ranked Pairs 0.161 0.162 0.164 0.165 0.182 0.166
Table 12: L1 Loss across preference distributions of networks trained using Impartial Culture preferences and the $T_{WT}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.289 0.29 0.355 0.388 0.537 0.379
Borda 0.934 0.935 0.737 0.666 0.935 0.611
Copeland 0.087 0.092 0.138 0.148 0.0 0.166
Schulze 0.735 0.744 0.355 0.338 0.287 0.281
SK 0.79 0.794 0.378 0.355 0.013 0.27
IRV 0.118 0.114 0.098 0.092 0.0 0.092
Black’s Rule 0.088 0.092 0.086 0.081 0.0 0.073
Ranked Pairs 0.102 0.097 0.097 0.091 0.0 0.085
Table 13: Test Accuracy across preference distributions of networks trained using Impartial Culture preferences and the $T_{WT}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.001 0.001 0.018 0.012 0.031 0.014
Borda 0.028 0.03 0.077 0.093 0.135 0.109
Copeland 0.0 0.0 0.001 0.001 0.0 0.002
Schulze 0.017 0.016 0.029 0.031 0.018 0.042
SK 0.017 0.016 0.028 0.027 0.02 0.039
IRV 0.163 0.162 0.163 0.164 0.182 0.162
Black’s Rule 0.022 0.023 0.066 0.085 0.018 0.103
Ranked Pairs 0.06 0.06 0.061 0.06 0.001 0.063
Table 14: L1 Loss across preference distributions of networks trained using Impartial Culture preferences and the $T_{CO}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.997 0.997 0.9 0.942 0.847 0.911
Borda 0.844 0.834 0.577 0.491 0.263 0.4
Copeland 1.0 1.0 0.999 0.998 1.0 0.977
Schulze 0.914 0.919 0.881 0.884 0.946 0.821
SK 0.913 0.917 0.888 0.914 0.894 0.829
IRV 0.106 0.109 0.101 0.1 0.0 0.111
Black’s Rule 0.88 0.873 0.638 0.534 0.9 0.433
Ranked Pairs 0.67 0.672 0.667 0.67 0.996 0.653
Table 15: Test accuracy across preference distributions of networks trained using Impartial Culture preferences and the $T_{CO}$ embedding.
Appendix F Participation Fine-Tuning
In this section we show the test loss for each of our rule-embedding pairs before and after fine-tuning (Table 16, Table 17, Table 18). Loss is calculated on a set of 1056 profiles with 44 voters. Across $T_{RF}$ and $T_{T}$ embeddings most rules experience only a minor increase in Rule Loss in exchange for significant decrease in Participation Loss during fine-tuning. Curiously, the $T_{WT}$ embedding is an exception to this; our fine-tuning process appears to optimize heavily for minimizing Participation Loss while greatly increasing Rule Loss. This suggests to us that the Weighted Tournament embedding provide information particularly well-suited to Participation Loss while being too complex for learning many rules. We also plot training loss during fine-tuning for each rule-embedding pair. Each plot shows 40 epochs of fine-tuning with a separate series for Rule Loss (L1 Loss), Participation Loss, and the sum of both loss terms. While Rule Loss occasionally increases a moderate amount during fine-tuning we see that combined loss consistently drops and in almost all cases the increase to Rule Loss is quite mild.
Before FT After FT
Target Rule Rule Part. Rule Part.
Plurality 0.0 0.247 0.095 0.005
Borda 0.165 0.0 0.165 0.0
Copeland 0.084 0.274 0.168 0.0
Schulze 0.064 0.352 0.076 0.069
Simpson-Kramer 0.068 0.313 0.074 0.097
IRV 0.158 0.0 0.158 0.0
Black’s Rule 0.037 0.291 0.161 0.001
Ranked Pairs 0.159 0.0 0.159 0.0
Table 16: Loss of our $T_{RF}$ models before and after fine-tuning with Participation Loss using $(m=11,n=44)$.
Before FT After FT
Target Rule Rule Part. Rule Part.
Plurality 0.098 0.528 0.113 0.16
Borda 0.034 0.352 0.042 0.226
Copeland 0.0 0.261 0.023 0.144
Schulze 0.017 0.361 0.036 0.132
Simpson-Kramer 0.02 0.358 0.037 0.139
IRV 0.099 0.27 0.102 0.184
Black’s Rule 0.027 0.339 0.037 0.211
Ranked Pairs 0.044 0.305 0.057 0.231
Table 17: Loss of our $T_{T}$ models before and after fine-tuning with Participation Loss using $(m=11,n=44)$.
Before FT After FT
Target Rule Rule Part. Rule Part.
Plurality 0.102 0.555 0.164 0.013
Borda 0.009 0.296 0.162 0.0
Copeland 0.163 0.0 0.162 0.0
Schulze 0.065 0.276 0.165 0.0
Simpson-Kramer 0.057 0.299 0.163 0.0
IRV 0.158 0.0 0.158 0.0
Black’s Rule 0.165 0.0 0.165 0.0
Ranked Pairs 0.159 0.0 0.159 0.0
Table 18: Loss of our $T_{WT}$ models before and after fine-tuning with Participation Loss using $(m=11,n=44)$.
Figure 4: Training Loss during fine-tuning for each rule using the $T_{RF}$ embedding.
Figure 5: Training Loss during fine-tuning for each rule using the $T_{T}$ embedding.
Figure 6: Training Loss during fine-tuning for each rule using the $T_{WT}$ embedding.
Figure 7: Training Loss during fine-tuning for each rule using the $T_{CO}$ embedding.